OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −one thirdx − 6 y = −one thirdx − 14 y = 3x − 6 y = 3x − 14

OpenStudy (shamil98):

Parallel lines have equal slopes. Put your equation in slope-intercept form first.

OpenStudy (shamil98):

After that, plug in (4,-2) and your slope (m) into y = mx + b and solve for b to find your y intercept of the equation.

OpenStudy (haseeb96):

since, y=mx+b....eq#1 y=3x+3 m=3 and b= 3 in parallel's condition slope remains same that is m =3 and (x,y)=(4,-2) -2=3(4)+b -2-12=b b=-14 now put it in eq#1 y=3x-14 it means that D is the correct answer